# AuToMATo: An Out-Of-The-Box Persistence-Based Clustering Algorithm

### v2

@@ 1 Introduction
Clustering techniques play a central role in understanding and interpreting data in a variety of fields. The idea is to divide a heterogeneous group of objects into groups based on a notion of similarity. This similarity is often measured with a distance or a metric on a data set. There exist many different clustering techniques [And73, DHS00], including hierarchical, centroid-based and density-based techniques, as well as techniques arising from probabilistic generative models. Each of these methods is proficient at finding clusters of a particular nature. Many of the most commonly used clustering algorithms require a selection of a parameters, a process which poses a considerable challenge when applying clustering to real-world problems.
@@
In this work, we present and implement AuToMATo (Automated Topological Mode Analysis Tool), a novel clustering algorithm based on the topological clustering algorithm ToMATo [CGOS13]. The latter summarizes the prominences of peaks of a density function in a so-called persistence diagram. The user then selects a prominence threshold $\tau$ and retains all peaks whose prominence is above this threshold, which results in the final clustering. A simple heuristic to select $\tau$ is to sort the peaks by decreasing prominence, and to look for the largest gap between two consecutive prominence values [CGOS13]. While yielding reasonable results in general, this procedure is not very robust to small changes in the prominence values.
@@
A more robust and sophisticated method is to perform a bottleneck bootstrap on the persistence diagram produced by ToMATo, which is precisely what AuToMATo does. That is, given a persistence diagram obtained by running ToMATo on a point cloud, AuToMATo produces a confidence region for that diagram with respect to the bottleneck distance, which translates into a choice of $\tau$ that determines the final clustering. While AuToMATo is not parameter-free per se, we provide default choices that make it perform well across the board. Unless stated otherwise, AuToMATo will henceforth refer to our algorithm with its parameters set to these defaults. We experimentally analyze the clustering performance of AuToMATo and we find that it not only outperforms parameter-free clustering algorithms, but often also even the best choice of hyperparameters for many parametric clustering algorithms. Parameter-free algorithms building on ToMATo exist in the literature, for example, in [CSTL21] the final clustering is determined by fitting a curve to the values of prominence, and in [BTO24] significant values are separated from non-significant ones by adapting the process that produces the persistence diagrams. Indeed, the former algorithm is one of those that AuToMATo is shown to outperform.
@@
We envision one important application of AuToMATo to be to the Mapper algorithm, introduced in [SMC+07]. Mapper constructs a graph that captures the topological structure of a data set. It relies on many parameters, one them being a clustering algorithm applied to various chunks of the data. Algorithms that depend heavily on a a good choice of a tunable hyperparameter are generally not good candidates for usage with Mapper, as the best choice for the hyperparameter can vary significantly over the different chunks, and manually choosing a different hyperparameter for each may not be possible in practice. Thus, most choices of hyperparameter will generally perform badly on some of the subsets, leading to undesired results of Mapper. Thus AuToMATo can be seen as progress towards finding optimal parameters for Mapper, which is active area of research [CMO18, CZW21, RHW23]. Running examples for Mapper with AuToMATo, we see that it is indeed a good choice for a clustering algorithm in this application when compared to parametric clustering algorithm such as DBSCAN.
@@ 2 Background
2.1 Persistence
Both ToMATo and AuToMATo rely on the theory of persistence [ELZ02, ZC05, Car14] to estimate the prominence of density peaks and to build the hierarchy of peaks. The first step towards using this theory is to construct a filtration from the space $X$ equipped with a density function $f$.
@@ 2.1 Persistence
Let $X$ be a topological space and let $f\colon X\to\mathbb{R}$ be a continuous function. The superlevelset filtration of $(X,f)$ is a family of superlevelsets $X_{\geq t}=f^{-1}\left([t,\infty)\right)$ together with the inclusion maps $X_{\geq t}\xhookrightarrow{}X_{\geq s}$ for $t\geq s$.
Our goal is to track the evolution of connected components of $X_{\geq t}$ as $t$ ranges from +$\infty$+ ∞ to $\infty$- ∞. In algebraic topology, connectivity is measured via 0-dimensional homology groups $\textrm{H}_{0}$ [Hat02]. There exist higher dimensional homology groups as well, for example, $\textrm{H}_{1}(X)$ provides information about the loops in $X$, $\textrm{H}_{2}(X)$ about the voids in $X$, etc. Homology groups are in fact functors, and, in addition to assigning vector spaces to topological spaces, they also assign linear maps to continuous maps between them.
Applying the $k$-dimensional homology functor (with coefficients in some field ${\mathbf{k}}$) to the superlevelset filtration yields a linear map
$\textrm{H}_{k}(i_{t}^{s}):\textrm{H}_{k}(X_{\geq t})\to\textrm{H}_{k}(X_{\geq s})$
for each $s,t$ with $t\geq s$ and for any $k\geq 0$. The collection of $\{\textrm{H}_{k}(X_{\geq t})\}_{t\in\mathbb{R}}$ together with induced linear maps is an example a persistence module [Car14]. In certain cases a persistence module can be expressed as a direct sum of ‘interval modules’, which can be thought of as the atomic building blocks of the theory.
@@
For an interval $[b,d)$ we denote by $\mathbb{I}_{[b,d)}$ the persistence module
$(\mathbb{I}_{[b,d)})_{t}=\begin{cases}{\mathbf{k}}&\textrm{ for }t\in[b,d)\\ 0&\textrm{ otherwise}\end{cases}\textrm{ and }\phi_{t}^{s}=\begin{cases}\mathrm{id}_{{\mathbf{k}}}&\textrm{ for }s\leq t,\textrm{ and }s,t\in[b,d)\\ 0&\textrm{ otherwise}\end{cases}.$
Each interval summand $\mathbb{I}_{[b,d)}$ represents a topological feature that is born at $t=b$ and dies at $s=d$. It can happen that there are features that never die, i.e. that $d=-\infty$. For example, when observing the superlevelset filtration of $(X,f)$ we have at least one connected component that persists to $\infty$- ∞ (provided $X$ is non-empty).
Definition 2.3.
@@
Let $\{(b_{l},d_{l})\}_{l}$ denote the birth and death times of connected components of the superlevelset filtration $\left\{X_{\geq t}\right\}_{t\in\mathbb{R}}$ associated to the density $f\colon X\to\mathbb{R}$. The associated persistence diagram, denoted by $\mathrm{Dgm}(X,f)$, is the multiset in the extended plane ¯2:={±}assignsuperscript¯$\overline{\mathbb{R}}^{2}:=\mathbb{R}\cup\{\pm\infty\}$ consisting of the points $\left\{(b_{l},d_{l})\right\}_{l}\subset\overline{\mathbb{R}}^{2}$ (counted with multiplicity) and the diagonal $\Delta:=\left\{(x,x)\mid x\in\overline{\mathbb{R}}\right\}$ (where each point on $\Delta$ has infinite multiplicity). For a given local maximum of $f$ with birth time $b_{l}$ and death time $d_{l}$ , we refer to the difference $d_{l}-b_{l}$ as its prominence or lifetime.
@@
The reason for working in the extended plane is that, provided that $f$ has a global maximum, the superlevelset filtration $X_{\geq t}$ will have a connected component that never dies, i.e., has death time equal to $\infty$- ∞. See the red graph in Figure 2 for an illustration.
@@
In the case we are interested in, when $X$ is a topological space and $f\colon X\to\mathbb{R}_{\geq 0}$ a density function, the persistence diagram $\mathrm{Dgm}(X,f)$ provides a summary of $f$: The points of $\mathrm{Dgm}(X,f)$ are in one-to-one correspondence with the local maxima of $f$, and twice the $L^{\infty}$-distance of a point to the diagonal $\Delta$ (i.e., its Euclidean vertical distance) equals its prominence.
Example 2.4.
Consider the function $f\colon\mathbb{R}\to\mathbb{R}$ in Figure 1. The function has two peaks. The first peak appears at $t=q$, and the second at $t=p$. At $t=s$, the connected component corresponding to $p$ is merged with that corresponding to $q$ since the peak $q$ is higher. So to $p$ we associate the point $(p,s)$ in Dgm($X,f$) and to $q$ we associate $(q,-\infty)$ (see Figure 1).
Figure 1: (a) A function $f\colon\mathbb{R}\to\mathbb{R}$ with local peaks $p$ and $q$. (b) Dgm($X,f$) for the function on the left.
A standard measure of similarity between persistence diagrams is the bottleneck distance [EH10, COGDS16].
Definition 2.5.
Let $\mathrm{Dgm}_{1}$ and $\mathrm{Dgm}_{2}$ be two persistence diagrams that have finitely many points off the diagonal. Let $\pi$ denote the set of bijections $\nu\colon\mathrm{Dgm}_{1}\to\mathrm{Dgm}_{2}$. Given points $x=(x_{1},x_{2})$ and $y=(y_{1},y_{2})$ in ¯$\overline{\mathbb{R}}^{2}$, let $\|x-y\|_{\infty}=\max\{|x_{1}-y_{1}|,|x_{2}-y_{2}|\}$ denote their $L^{\infty}$-distance, where we set $(+\infty)-(+\infty)=(-\infty)-(-\infty)=0$. Then, the bottleneck distance between $\mathrm{Dgm}_{1}$ and $\mathrm{Dgm}_{2}$ is defined as
Persistence diagrams are stable with respect to small perturbations [CSEH07], i.e. for ${f,g\colon X\to\mathbb{R}}$
$W_{\infty}(\textrm{Dgm}(X,f),\textrm{Dgm}(X,g))\leq\|f-g\|_{\infty}$
under certain assumptions on $(X,f)$ and $(X,g)$. This holds, for example, if $f,g$ are Morse functions on compact manifolds, piecewise linear functions on simplicial complexes, etc. There are versions of the stability result that hold more generally, for example [CdSO14, CdSMS16, CCSG+09].
Example 2.6.
Consider a function $f$ defined on a compact subset of $\mathbb{R}$ and a piecewise linear approximation of $f$, $g$ (Figure 2). Since the functions are close to each other with respect to the $L^{\infty}$ distance, the stability theorem guarantees that the corresponding persistence diagrams are close with respect to the bottleneck distance. Indeed, the number of prominent peaks is the same in both cases.
Figure 2: (a) A function $f\colon K\subset\mathbb{R}\to\mathbb{R}$ and an approximation of $f$, $g$. (b) Dgm($X,f$) consists of red dots, while Dgm($X,g$) consists of the blue dots.
2.2 ToMATo: a Persistence-Based Clustering Algorithm
@@
We now briefly recall how the ToMATo clustering algorithm works. Given a point cloud $X$ ToMATo relies on the assumption that the points of $X$ were sampled according to some unknown density function $f$. In a nutshell, ToMATo infers information about the local maxima of $f$ by applying the above procedure to an estimate of $f$. ToMATo takes as input:
@@
  • A neighborhood graph $G$ on the points of $X$. Chazal et al. mostly use the $\delta$-Rips graph and the $k$-nearest neighbor graph. [^1] Given a point cloud, both of these undirected graphs have the set of data points as their vertex set. In the case of the $\delta$-Rips graph, two vertices are connected iff they are at a distance of at most $\delta$ apart, whereas in the $k$-nearest neighbor graph, a data point is connected to another iff the latter is among the $k$-nearest neighbors of the first.
@@
  • A density estimator $\hat{f}$. Each vertex $v$ of $G$ is assigned a non-negative value $\hat{f}(v)$ that corresponds to the estimated density at $v$. Chazal et al. propose two possible density estimators: the truncated Gaussian kernel density estimator and the distance-to-measure density, originally introduced in [BCCS+11]. [^2] For a smoothing parameter $m\in(0,1)$, and a given data point $x$, its empirical (unnormalized) distance-to-measure density is given by $\hat{f}(x)=\left(\frac{1}{k}\sum_{y\in N_{k}(x)}\|x-y\|^{2}\right)^{-\frac{1}{2}},$ where $k=\lceil mn\rceil$, $N_{k}(x)$ denotes the set of the $k$ nearest neighbors of $x$, and $n$ is the cardinality of the data set.
@@
    Apply a hill-climbing algorithm on $G$. Construct the neighborhood graph $G$ on the points of $X$, and construct a directed subgraph $\Gamma$ of $G$ as follows: at each vertex $v$ of $G$, place a directed edge from $v$ to its neighbor with highest value of $\hat{f}$, provided that that value is higher than $\hat{f}(v)$. If all neighbors of $v$ have lower values, $v$ is a peak of $\hat{f}$. This yields a collection of directed edges that form a spanning forest of the graph $G$, consisting of one tree for each local maximum of $\hat{f}$. In particular, these trees yield a partition of the elements of $X$ into pairwise disjoint sets that serves as a candidate clustering on $X$.
@@
    Construct the persistence diagram. Construct the persistence diagram $\mathrm{Dgm}(G,\hat{f})$ associated to the superlevelset filtration of $\hat{f}\colon G\to\mathbb{R}$.
@@
    Merge non-significant clusters. Iteratively merge every cluster of prominence less than $\tau$ of the candidate clustering found in Step 2 into its parent cluster, i.e., into the cluster corresponding to the local maximum that it gets merged into in the superlevelset filtration of $\hat{f}\colon G\to\mathbb{R}$. ToMATo outputs the resulting clustering of points of $X$, in which every cluster has prominence at least $\tau$ by construction.
The reason why we can expect the persistence diagram of the approximated density to be “close” to the original one stems from the stability of persistence diagrams under the bottleneck distance, as explained in Section 2.1 and illustrated in Figure 2.
In practice, the user must run ToMATo twice. First, ToMATo is run with $\tau=+\infty$, which is equivalent to computing the birth and death time of each local maximum of $\hat{f}$ and hence the persistence diagram $\mathrm{Dgm}(G,\hat{f})$. From the diagram $\mathrm{Dgm}(G,\hat{f})$ the user then determines a merging parameter $\tau$ by visually identifying a large gap in $\mathrm{Dgm}(G,\hat{f})$ separating, say, $C$ points corresponding to highly prominent peaks from the rest of the points. Then, ToMATo is run a second time with $\tau$ set to that value, which results in the final clustering of $X$ into $C$ clusters.
@@ 2.3 The Bottleneck Bootstrap
We now review the second ingredient of our clustering scheme, a procedure known as the bottleneck bootstrap. Introduced in [CFL+17, Section 6], it is used to separate significant features in persistence diagrams from non-significant ones. While it may be used in more general settings, we will restrict ourselves to the scenario of Section 2.2
Suppose that $X$ is a sample consisting of $n$ data points, drawn according to some unknown probability density function $f\colon K\to[0,1]$, $K\subset\mathbb{R}^{n}$, and let $\mathcal{D}:=\mathrm{Dgm}(K,f)$ denote the corresponding (unknown) persistence diagram (we assume the density $f$ and all of its estimates to be normalized here for ease of exposition). We estimate $f$ and the connectivity of $K$ with a density estimator and a neighborhood graph, respectively (as explained in Section 2.2). This allows us to compute $\widehat{\mathcal{D}}:=\mathrm{Dgm}(X,\hat{f})$, where $\hat{f}$ is the estimate of $f$). The diagram $\widehat{\mathcal{D}}$ serves as an estimate for $\mathcal{D}$.
Given a confidence level $\alpha\in(0,1)$ and a number $B\in\mathbb{Z}_{\geq 1}$ of bootstrap iterations, the bottleneck bootstrap gives an estimate $\widehat{q}_{\alpha}$ of $q_{\alpha}$, which is defined by
$\mathbb{P}(\sqrt{n}W_{\infty}(\mathcal{D},\widehat{\mathcal{D}})\leq q_{\alpha})=1-\alpha.$ (1)
In order to produce $\widehat{q}_{\alpha}$, we first approximate $f$ with the empirical measure $P_{n}$ on $X$ that assigns the probability mass of $1/n$ to each data point (note that this generally does not coincide with $\hat{f}$). This allows us to estimate the distribution
$F(z):=\mathbb{P}(\sqrt{n}W_{\infty}(\mathcal{D},\widehat{\mathcal{D}})\leq z)$
@@
$\widehat{F}(z):=\mathbb{P}(\sqrt{n}W_{\infty}(\widehat{\mathcal{D}}^{*},\widehat{\mathcal{D}})\leq z),$
where $\widehat{\mathcal{D}}^{*}:=\mathrm{Dgm}(X^{*},\hat{f}^{*})$ is the persistence diagram corresponding to a sample $X^{*}$ of size $n$ drawn from $P_{n}$, and the density $\hat{f}^{*}$ and the connectivity of $X^{*}$ are estimated using the same estimators as before. Note that $X^{*}$ may be thought of as a sample drawn from $X$ with replacement. The distribution $\widehat{F}$ itself is approximated by Monte Carlo as follows. We draw $B$ samples $X_{1}^{*},\dots,X_{B}^{*}$ of size $n$ from $P_{n}$, and for each of these $B$ samples, we compute the persistence diagram $\widehat{\mathcal{D}}_{i}^{*}:=\mathrm{Dgm}(X_{i}^{*},\hat{f}_{i}^{*})$ and the quantity $T_{i}^{*}:=\sqrt{n}W_{\infty}(\widehat{\mathcal{D}}_{i}^{*},\widehat{\mathcal{D}})$, $i=1,\dots,B$. Finally, we use the function
$\widetilde{F}(z):=\frac{1}{B}\sum_{i=1}^{B}\mathbf{1}_{[0,z]}(T_{i}^{*})$

### v3

@@ 1 Introduction
Clustering techniques play a central role in understanding and interpreting data in a variety of fields. The idea is to divide a heterogeneous group of objects into groups based on a notion of similarity. This similarity is often measured with a distance or a metric on a data set. There exist many different clustering techniques (Anderberg, 1973; Duda et al., 2000), including hierarchical, centroid-based and density-based techniques, as well as techniques arising from probabilistic generative models. Each of these methods is proficient at finding clusters of a particular nature. Many of the most commonly used clustering algorithms require a selection of parameters, a process which poses a considerable challenge when applying clustering to real-world problems.
@@
In this work, we present and implement AuToMATo (Automated Topological Mode Analysis Tool), a novel clustering algorithm based on the topological clustering algorithm ToMATo (Chazal et al., 2013). The latter summarizes the prominences of peaks of a density function in a so-called persistence diagram. The user then selects a prominence threshold $\tau$ and retains all peaks whose prominence is above this threshold, which results in the final clustering. A simple heuristic to select $\tau$ is to sort the peaks by decreasing prominence, and to look for the largest gap between two consecutive prominence values (Chazal et al., 2013). While yielding reasonable results in general, this procedure is not very robust to small changes in the prominence values.
@@
A more robust and sophisticated method is to perform a bottleneck bootstrap on the persistence diagram produced by ToMATo, which is precisely what AuToMATo does. That is, given a persistence diagram obtained by running ToMATo on a point cloud, AuToMATo produces a confidence region for that diagram with respect to the bottleneck distance, which translates into a choice of $\tau$ that determines the final clustering. While AuToMATo is not parameter-free per se, we provide default choices that make it perform well across the board. Unless stated otherwise, AuToMATo will henceforth refer to our algorithm with its parameters set to these defaults. We experimentally analyze the clustering performance of AuToMATo and we find that it not only outperforms parameter-free clustering algorithms, but often also even the best choice of hyperparameters for many parametric clustering algorithms. Parameter-free algorithms building on ToMATo exist in the literature, for example, in Cotsakis et al. (2021) the final clustering is determined by fitting a curve to the values of prominence, and in Bois et al. (2024) significant values are separated from non-significant ones by adapting the process that produces the persistence diagrams. Indeed, the former algorithm is one of those that AuToMATo is shown to outperform.
@@
We envision one important application of AuToMATo to be to the Mapper algorithm, introduced in Singh et al. (2007). Mapper constructs a graph that captures the topological structure of a data set. It relies on many parameters, one of them being a clustering algorithm applied to various chunks of the data. Algorithms that depend heavily on a good choice of a tunable hyperparameter are generally not good candidates for usage with Mapper, as the best choice for the hyperparameter can vary significantly over the different chunks, and manually choosing a different hyperparameter for each may not be possible in practice. Thus, most choices of hyperparameter will generally perform badly on some of the subsets, leading to undesired results of Mapper. Thus, AuToMATo can be seen as progress towards finding optimal parameters for Mapper, which is an active area of research (Carrière et al., 2018; Chalapathi et al., 2021; Rosen et al., 2023). Running examples for Mapper with AuToMATo, we see that it is indeed a good choice for a clustering algorithm in this application when compared to parametric clustering algorithm such as DBSCAN.
@@ 2 Background
2.1 Persistence and the ToMATo clustering algorithm
Both ToMATo and AuToMATo rely on the theory of persistence (Edelsbrunner et al., 2002; Zomorodian & Carlsson, 2005; Carlsson, 2014) to quantify the prominence of peaks of (an estimate of) a density function, and to build a hierarchy of peaks. Given a topological space $X$ equipped with a density function $f\colon X\to\mathbb{R}_{\geq 0}$, the first step of persistence is to build a filtration from $X$.
@@ 2.1 Persistence and the ToMATo clustering algorithm
Let $X$ be a topological space, and let $f\colon X\to\mathbb{R}$ be continuous. The superlevel set filtration of $(X,f)$ is the family of superlevel sets $\left\{X_{\geq t}\mid t\in\mathbb{R}\right\}$, where $X_{\geq t}\coloneqq f^{-1}\left([t,\infty)\right)$.
In the following, we assume for ease of exposition that all local extrema of $f$ have distinct values. The idea underlying ToMATo is to track the evolution of (the number of) connected components of $X_{\geq t}$ as $t$ ranges from $\infty$ to -$\infty$. In that process, the number of connected components of $X_{\geq t}$ remains constant, unless $t$ passes through the value of a local extremum of $f$. As $t$ passes through the value of a local maximum, a new connected component is “born” and added to the superlevel set $X_{\geq t}$. Similarly, as $t$ passes through the value of a local minimum, two connected components of $X_{\geq t}$ are merged into one. ToMATo builds a hierarchy of local maxima of $f$ by declaring that, as two components get merged, the component corresponding to the local maximum with higher value absorbs the other one and persists, whereas the component corresponding to the local maximum with lower value “dies”. Therefore, to each local maximum we associate a pair $(b,d)$ where $b$ denotes the birth and $d$ the death time, respectively. The evolution of the connected components can be concisely recorded in a persistence diagram.
@@
Let $\{(b_{l},d_{l})\}_{l}$ denote the birth and death times of connected components of the superlevel set filtration $\left\{X_{\geq t}\right\}_{t\in\mathbb{R}}$ associated to the density $f\colon X\to\mathbb{R}$. The associated persistence diagram, denoted by $\mathrm{Dgm}(X,f)$, is the multiset in the extended plane $\overline{\mathbb{R}}^{2}\coloneqq\mathbb{R}\cup\{\pm\infty\}$ consisting of the points $\left\{(b_{l},d_{l})\right\}_{l}\subset\overline{\mathbb{R}}^{2}$ (counted with multiplicity) and the diagonal $\Delta\coloneqq\left\{(x,x)\mid x\in\overline{\mathbb{R}}\right\}$ (where each point on $\Delta$ has infinite multiplicity). For a given local maximum of $f$ with birth time $b_{l}$ and death time $d_{l}$ , we refer to the difference $d_{l}-b_{l}$ as its prominence or lifetime.
@@
The reason for working in the extended plane is that, provided that $f$ has a global maximum, the superlevel set filtration $X_{\geq t}$ will have a connected component that never dies, that is, has death time equal to -$\infty$. See the red graph in Figure 1 for an illustration.
@@
The persistence diagram $\mathrm{Dgm}(X,f)$ provides a summary of $f$. The points of $\mathrm{Dgm}(X,f)$ are in one-to-one correspondence with the local maxima of $f$, and twice the $L^{\infty}$-distance of a point to the diagonal $\Delta$ (that is, its Euclidean vertical distance) equals its prominence.
@@
We now outline how the ToMATo clustering algorithm works. Given a point cloud $X$ ToMATo relies on the assumption that the points of $X$ were sampled according to some unknown density function $f$. In a nutshell, ToMATo infers information about the local maxima of $f$ by applying the above procedure to an estimate of $f$. ToMATo takes as input:
@@
  • A neighborhood graph $\mathcal{G}$ on the points of $X$. Chazal et al. mostly use the $\delta$-Rips graph and the $k$-nearest neighbor graph. [^1] Given a point cloud, both of these undirected graphs have the set of data points as their vertex set. In the case of the $\delta$-Rips graph, two vertices are connected iff they are at a distance of at most $\delta$ apart, whereas in the $k$-nearest neighbor graph, a data point is connected to another iff the latter is among the $k$-nearest neighbors of the first.
@@
  • A density estimator $\hat{f}$. Each vertex $v$ of $\mathcal{G}$ is assigned a non-negative value $\hat{f}(v)$ that corresponds to the estimated density at $v$. Chazal et al. propose two possible density estimators: the truncated Gaussian kernel density estimator and the distance-to-measure density, originally introduced in Biau et al. (2011). [^2] For a smoothing parameter $m\in(0,1)$, and a given data point $x$, its empirical (unnormalized) distance-to-measure density is given by $\hat{f}(x)=\left(\frac{1}{k}\sum_{y\in N_{k}(x)}\|x-y\|^{2}\right)^{-\frac{1}{2}},$ where $k=\lceil mn\rceil$, $N_{k}(x)$ denotes the set of the $k$ nearest neighbors of $x$, and $n$ is the cardinality of the data set.
@@
    Apply a hill-climbing algorithm on $\mathcal{G}$. Construct the neighborhood graph $\mathcal{G}$ on the points of $X$, and construct a directed subgraph $\mathcal{G}^{\prime}$ of $\mathcal{G}$ as follows: at each vertex $v$ of $\mathcal{G}$, place a directed edge from $v$ to its neighbor with highest value of $\hat{f}$, provided that that value is higher than $\hat{f}(v)$. If all neighbors of $v$ have lower values, $v$ is a peak of $\hat{f}$. This yields a collection of directed edges that form a spanning forest of the graph $\mathcal{G}$, consisting of one tree for each local maximum of $\hat{f}$. In particular, these trees yield a partition of the elements of $X$ into pairwise disjoint sets that serves as a candidate clustering on $X$.
@@
[…]
@@
    Merge non-significant clusters. Iteratively merge every cluster of prominence less than $\tau$ of the candidate clustering found in Step 2 into its parent cluster, that is, into the cluster corresponding to the local maximum that it gets merged into in the superlevel set filtration of $\hat{f}\colon\mathcal{G}\to\mathbb{R}$. ToMATo outputs the resulting clustering of points of $X$, in which every cluster has prominence at least $\tau$ by construction.
The reason why we can expect the persistence diagram of the approximated density to be “close” to the original one stems from the stability of persistence diagrams under the bottleneck distance (explained in Section 2.2). This is illustrated in Figure 1.
[…]
@@ 2.2 The bottleneck bootstrap
The bottleneck bootstrap, introduced in Chazal et al. (2017, Section 6), is used to separate significant features in persistence diagrams from non-significant ones. While it may be used in more general settings, we will restrict ourselves to the scenario of Section 2.1.
We first review the bottleneck distance, which is the standard distance measure between persistence diagrams (Edelsbrunner & Harer, 2010; Chazal et al., 2016).
Definition 2.3.
Let $\mathrm{Dgm}_{1}$ and $\mathrm{Dgm}_{2}$ be two persistence diagrams that have finitely many points off the diagonal. Let $\pi$ denote the set of bijections $\nu\colon\mathrm{Dgm}_{1}\to\mathrm{Dgm}_{2}$. Given points $x=(x_{1},x_{2})$ and $y=(y_{1},y_{2})$ in $\overline{\mathbb{R}}^{2}$, let $\|x-y\|_{\infty}=\max\{|x_{1}-y_{1}|,|x_{2}-y_{2}|\}$ denote their $L^{\infty}$-distance, where we set $(+\infty)-(+\infty)=(-\infty)-(-\infty)=0$. Then, the bottleneck distance between $\mathrm{Dgm}_{1}$ and $\mathrm{Dgm}_{2}$ is defined as
We now outline the bottleneck bootstrap. It relies on the following theorem, which summarizes the relevant results of Chazal et al. (2017, Section 6).
Theorem 2.4 (Chazal et al. (2017)).
Let $X\subseteq\mathbb{R}^{N}$ be a sample consisting of $n$ data points drawn according to a probability density function $f\colon K\to[0,1]$, $K\subset\mathbb{R}^{N}$. Denote by $\mathcal{D}\coloneqq\mathrm{Dgm}(K,f)$ and $\widehat{\mathcal{D}}\coloneqq\mathrm{Dgm}(X,f)$ the corresponding unknown and estimated, respectively, persistence diagrams of superlevel sets. Given a confidence level $\alpha\in(0,1)$, define $q_{\alpha}$ by
$\mathbb{P}(\sqrt{n}W_{\infty}(\mathcal{D},\widehat{\mathcal{D}})\leq q_{\alpha})=1-\alpha.$
Then a consistent estimator for $q_{\alpha}$ is given by $\widehat{q}_{\alpha}$, which in turn is defined by
$\mathbb{P}(\sqrt{n}W_{\infty}(\widehat{\mathcal{D}}^{*},\widehat{\mathcal{D}})\leq\widehat{q}_{\alpha})=1-\alpha.$
Here, $\widehat{\mathcal{D}}^{*}\coloneqq\mathrm{Dgm}(X^{*},f)$ denotes the random persistence diagram constructed from a sample $X^{*}$ of size $n$ drawn according to the empirical measure $P_{n}$ on $X$, where $P_{n}$ is defined as the probability measure on $X$ that assigns the probability mass $1/n$ to each data point in $X$.
In our setting, the theorem above is applied as follows. Given the sample $X\subseteq\mathbb{R}^{N}$, we estimate $f$ and the connectivity of $K$ with a density estimator and a neighborhood graph, respectively (as explained in Section 2.1). This allows us to compute $\widehat{\mathcal{D}}\coloneqq\mathrm{Dgm}(X,f)$, which, in turn, serves as an estimate of $\mathcal{D}$. The empirical measure $P_{n}$ on $X$ serves as an approximation of the unknown probability measure $f$, and using this, Theorem 2.4 allows us to approximate the distribution
$F(z)\coloneqq\mathbb{P}(\sqrt{n}W_{\infty}(\mathcal{D},\widehat{\mathcal{D}})\leq z)$
@@
$\widehat{F}(z)\coloneqq\mathbb{P}(\sqrt{n}W_{\infty}(\widehat{\mathcal{D}}^{*},\widehat{\mathcal{D}})\leq z),$
where $\widehat{\mathcal{D}}^{*}\coloneqq\mathrm{Dgm}(X^{*},f)$ is a random quantity. Note that $X^{*}$ may be thought of as a sample drawn from $X$ with replacement.
[…]
$\widetilde{F}(z)\coloneqq\frac{1}{B}\sum_{i=1}^{B}\mathbf{1}_{[0,z]}(T_{i}^{*})$